OpenStudy (vispla22):

what is sin(theta-90)

hartnn (hartnn):

use these 2 relations \(\Large \sin (90-\theta ) = \cos \theta \\ \Large \sin(-\theta ) = -\sin \theta \)

OpenStudy (vispla22):

could you say that sin(90-theta) is equal to sin(90)-sin(theta)?

hartnn (hartnn):

no, it doesn't work that way \(\sin (A+B) \ne \sin A+\sin B \\\Large \sin (A+B) = \sin A \cos B +\cos A \sin B\) you can use this formula too

OpenStudy (vispla22):

could you explain it to me???? i still don't really get it

hartnn (hartnn):

sure, see, we have the formula sin (90-theta) = cos theta but we need sin (theta - 90) so we do \(\huge \sin (90-\theta) = \sin [-(90-\theta )]\) got this step ?

hartnn (hartnn):

i mean \(\huge \sin (\theta-90) = \sin [-(90-\theta )]\)

OpenStudy (vispla22):

yeah

hartnn (hartnn):

good, so now we use \( \sin (-x)=-\sin x \) so we get \(\Large \sin [-(90-\theta)] = - \sin (90-\theta) = -\cos \theta\) makes sense ?

OpenStudy (vispla22):

oooooooohhhhh

OpenStudy (vispla22):

thx!

OpenStudy (vispla22):

so it could also work the other way around and say cos(theta-90)=-sintheta?

hartnn (hartnn):

but \(\Large \cos (-x) = \cos x\) hence \(\cos (90-\theta ) = \cos (\theta -90) = \sin \theta\) :)

OpenStudy (vispla22):

k thx again!

hartnn (hartnn):

welcome ^_^ you seem to be new here , \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)